# Energetically consistent localised APE budgets for local and regional studies of stratified flow energetics

### v1

@@ 4.4 Remarks on the eddy (turbulent) kinetic energy budget
Since skew-diffusion associated with the meso-scale eddy parameterisation acts as a net source of eddy KE rather than of eddy APE, it is useful to conclude this section with some comments on the eddy KE budget and how the KE and APE cascades may help constrain mixing parameterisations. briefly comment on the eddy KE budget. Under stationary and homogeneous conditions, the main sources of eddy KE are contributed by:
@@
    the downscale energy transfer associated with the mixing of momentum. To be specific, we assume that it is primarily contributed by the vertical shear
@@
Summing up all these contributions leads to the following balance
@@
To show the dependence of this balance on the Richardson number, it is useful to divide this relation by $\overline{N}^{2}$, which yields
@@
is the Richardson number. This relation shows that the different turbulent mixing parameters $K_{v}$, $\kappa_{gm}$, and $K_{\rho}$ are not independent from each other. Note that by using the relation $K_{\rho}\approx\Gamma\varepsilon_{k}^{t}/\overline{N}^{2}$, the above relation can be written in the form
@@
and can be interpreted as a diagnostic energy balance being potentially useful to predict the turbulent kinetic energy dissipation rate, which is a crucial elements of energetically consistent meso-scale eddy parameterisations such as in the GEOMETRIC framework e.g. Marshall et al. (2012); Mak et al. (2018); Torres et al. (2023) or in Jansen et al. (2015, 2019) for instance. In regions where the vertical shear can be predicted by thermal wind balance
@@
the TKE budget may also be written in the form
@@
Eq. (100) shows that in this regime, the APE cascade converting mean APE into eddy APE behaves analogously to the KE cascade converting mean KE into eddy KE, similarly as Lindborg (2006)’s relations (3), provided that
@@
which appears to be closely related to the case discussed by Greatbatch and Lamb (1990). Determining to what extent the present results can inform the further developments of such a parameterisation is beyond the scope of this paper and will be addressed in a subsequent study.
@@ 5 Summary and discussion
In this study, we have revisited the mean/eddy decomposition theory for local Available Potential Energy (APE) density, focusing on its application to the characterization of meso-scale eddies, small-scale irreversible mixing, and their interaction with the large-scale circulation. While irreversible mixing and eddy features occur at scales much smaller than the planetary-scale circulation, these processes play a crucial role in the global energy budget, necessitating careful consideration to ensure that forcing and dissipation terms are computed in a mutually consistent manner. Our approach addresses the consistency issues associated with the definition of localized forms of local APE density, which are often used to describe the energetics of these processes. Our analysis demonstrates that a consistent formulation should be based on partitioning the local APE density into mean and eddy components. The resulting eddy APE density measures forces against the local mean density profile, aligning with physical intuition. This formulation differs from heuristic localized forms of APE density, which ignore the dependence on the Lorenz reference state. Our exact eddy APE density retains some dependence on the Lorenz reference state, differing from the heuristic form by a factor of $\Lambda=\overline{N}^{2}/N_{0}^{2}$. Previous studies indicate that while fluid parcels are very close to their reference position in most of the ocean interior, where $\Lambda\approx 1$, this is generally not the case in the polar regions or the Gulf Stream area (e.g., Saenz et al. (2015); Tailleux (2016, 2013a)) where ou results should be of most practical use.

### v2

@@ 4.4 Remarks on the eddy (turbulent) kinetic energy budget
Since skew diffusion associated with the meso-scale eddy parameterisation acts as a net source of eddy KE rather than of eddy APE, it is useful to conclude this section with some comments on the eddy KE budget and how the KE and APE cascades may help constrain mixing parameterisations. Under stationary and homogeneous conditions, the main sources of eddy KE are:
@@
    the downscale energy transfer associated with the mixing of momentum, primarily contributed by the vertical shear,
@@
Summing up all these contributions leads to the following balance:
@@
To show the dependence of this balance on the Richardson number, it is useful to divide this relation by $\overline{N}^{2}$, yielding
@@
is the Richardson number. This relation shows that the different turbulent mixing parameters $K_{v}$, $\kappa_{gm}$, and $K_{\rho}$ are not independent from each other. Notably, by using the relation $K_{\rho}\approx\Gamma\varepsilon_{k}^{t}/\overline{N}^{2}$, the above relation can be written as
@@
and can be interpreted as a diagnostic energy balance potentially useful for predicting the turbulent kinetic energy dissipation rate, which is a crucial element of energetically consistent meso-scale eddy parameterisations such as in the GEOMETRIC framework (e.g., Marshall et al. (2012); Mak et al. (2018); Torres et al. (2023)) or in Jansen et al. (2015, 2019), for instance. In regions where the vertical shear can be predicted by thermal wind balance,
@@
the TKE budget may also be written as
@@
Eq. (100) shows that in this regime, the APE cascade converting mean APE into eddy APE behaves analogously to the KE cascade converting mean KE into eddy KE, similarly to Lindborg (2006)’s relations (3), provided that
@@
which appears to be closely related to the case discussed by Greatbatch and Lamb (1990). Determining to what extent the present results can inform further developments of such parameterisations is beyond the scope of this paper and will be addressed in a subsequent study.
@@ 5 Summary and discussion
In this study, we have revisited the mean/eddy decomposition theory for local Available Potential Energy (APE) density, focusing on its application to the characterization of meso-scale eddies, small-scale irreversible mixing, and their interaction with the large-scale circulation. While irreversible mixing and eddy features occur at scales much smaller than the planetary-scale circulation, these processes play a crucial role in the global energy budget, necessitating careful consideration to ensure that forcing and dissipation terms are computed in a mutually consistent manner. Our approach addresses the consistency issues associated with the definition of localized forms of local APE density, which are often used to describe the energetics of these processes. Our analysis demonstrates that a consistent formulation should be based on partitioning the local APE density into mean and eddy components. The resulting eddy APE density measures forces against the local mean density profile, aligning with physical intuition. This formulation differs from heuristic localized forms of APE density, which ignore the dependence on the Lorenz reference state. Our exact eddy APE density retains some dependence on the Lorenz reference state, differing from the heuristic form by a factor of $\Lambda=\overline{N}^{2}/N_{0}^{2}$. Previous studies indicate that while fluid parcels are very close to their reference position in most of the ocean interior, where $\Lambda\approx 1$, this is generally not the case in the polar regions or the Gulf Stream area (e.g., Saenz et al. (2015); Tailleux (2016, 2013a)) where our results should be of most practical use.